Determine whether the equation is exact. If it is, then solve it. 3 et(3y – t)dt+ (7+9e) dy = 0 Select the correct choice belThank you!!
Answer #1

Yoga on We have, ze (3y -t) et + (7 +966) dy = 0 Now , Mly, t) = 3et (3y-t) and Nly, t) = 7+ get qet Then, DN at aM = get ay

Please feel free to ask any doubts regarding the solution and please rate positively.

Kind regards.